# Conformal Prediction and MLLM aided Uncertainty Quantification in Scene Graph Generation

### v1

@@ 5.3 Comparative Uncertainty Quantification of SGG Methods
The empirical coverage of the object, predicate, and triplet prediction sets, obtained by PC-SGG for each SGG method is shown in Table 1. For all the methods, the object prediction sets fail to achieve the class-conditional coverage guarantee of $95\%$ (Eq 2). In comparison, the predicate prediction sets of VCTREE, SQUAT, and BGNN come close to the coverage guarantee of $90\%$. However, for all the methods, $AvgSize$ of the predicate prediction sets is significantly higher than that of the objects. This highlights the uncertainty associated with predicate classification is significantly higher than object classification. As such, the predicate prediction sets larger to meet coverage guarantees. Overall, SQUAT and BGNN achieve the best empirical coverage for their triplet prediction sets, with SQUAT having the smallest $AvgSize$ for its predicate prediction sets. From the $CovGap$ values, it can also be observed that the predicate prediction sets of SQUAT achieve the lowest class-wise deviation from the class-conditional coverage guarantee. For the object prediction sets, BGNN has lowest $CovGap$. The empirical coverage values of the triplet prediction set, $Cov_{T}$, also verify Theorem 1 (check supplementary) since it is approximately close to the product of the object and predicate prediction sets’ $Cov$ values. However, the $Cov_{T}$ values of none of the models achieve the marginal coverage guarantee of $(1-\alpha_{o})(1-\alpha_{r})$. This observation not only highlights the high predictive uncertainty of existing SGG methods but also sheds light on possible distribution shifts in the train and test splits of SGG datasets that make computing well-calibration conformal quantiles very challenging.
@@ 5.4 Results of MLLM-based Post-Processing
Table 3: Comparison of Recall@K(R@K) and mean-Recall@K(mR@K) with coverage-Recall@K(cR@K) and coverage-mean-Recall@K(cmR@K) when PC-SGG is added to each method. Bold values show maximum improvement.
@@ 5.5 Comparsion of SGG Metrics after incorporating PC-SGG
Table 3 shows the effectiveness of adding PC-SGG to the different SGG methods by comparing standard Recall@K (R@K) and mean-Recall@K (mR@K) with their conformal prediction-based counterparts, coverage-Recall@K (cR@K) and coverage-mean-Recall@K (cmR@K)5.2. When PC-SGG is added, the recall-hit rate improves significantly since the set-valued triplet predictions provide a much greater chance of covering the actual ground truth than the standard SGG scenario, where each detected triplet has only one prediction as opposed to a set of predictions. On average, across all the SGG methods, cR@$50$ improves by $15.70$ percentage points over R@$50$.
Additionally, since PC-SGG is designed to achieve class-conditional coverage guarantees, adding it to any SGG method significantly improves performance on the tail classes, as evidenced by the massive performance gains when comparing mR@K with cmR@K. On average, across all the SGG methods, cmR@$50$ improves by $21.57$ percentage points over mR@$50$. Note that the R@K and mR@k values of the standalone models are different from the ones reported in their respective papers since we retrain them on a new training split of VG150. Some qualitative results are provided in the supplementary.
@@ 6 Conclusion
Acknowledgments.This work was partially supported by NSF grants OAC-2411453, CCF-2008020 and DURIP N000141812252.

### v2

@@ 5.3 Comparative Uncertainty Quantification of SGG Methods
The empirical coverage of the object, predicate, and triplet prediction sets, obtained by PC-SGG for each SGG method is shown in Table 1. For all the methods, the object prediction sets fail to achieve the class-conditional coverage guarantee of $95\%$ (Eq 2). In comparison, the predicate prediction sets of VCTREE, SQUAT, and BGNN come close to the coverage guarantee of $90\%$. However, for all the methods, $AvgSize$ of the predicate prediction sets is significantly higher than that of the objects. This shows that predicate classification uncertainty is significantly higher than object classification. As such, the predicate prediction sets are larger to meet coverage guarantees. Overall, SQUAT and BGNN achieve the best empirical coverage for their triplet prediction sets, with SQUAT having the smallest $AvgSize$ for its predicate prediction sets. From the $CovGap$ values, it can also be observed that the predicate prediction sets of SQUAT achieve the lowest class-wise deviation from the class-conditional coverage guarantee. For the object prediction sets, BGNN has lowest $CovGap$. The empirical coverage values of the triplet prediction set, $Cov_{T}$, also verify Theorem 1 (check supplementary) since it is approximately close to the product of the object and predicate prediction sets’ $Cov$ values. However, the $Cov_{T}$ values of none of the models achieve the marginal coverage guarantee of $(1-\alpha_{o})(1-\alpha_{r})$ shown in Corollary 1. This observation highlights the high predictive uncertainty of existing SGG methods, as well as, shed light on possible distribution shifts in the train and test splits of SGG datasets that make computing well-calibration conformal quantiles very challenging under exchangeability assumptions.
@@ 5.4 Results of MLLM-based Post-Processing
Table 3: Comparison of cR@K and cmR@K of the different SGG methods when PC-SGG is added. Performance of these methods for standard SGG based on R@K and mR@K are also shown.
@@ 5.5 Comparsion of SGG Metrics after incorporating PC-SGG
[…]
@@ 6 Conclusion
Acknowledgments.This work was partially supported by NSF grants OAC-2411453, CCF-2008020, CNS-2312395, CMMI-2326309 and DURIP N000141812252.